# GIM: A Million-scale Benchmark for Generative Image Manipulation Detection and Localization

### v1

@@ 6.1 GIM Dataset Configuration
The specific quantities and details of the dataset are presented in Table 9, GIM has a total number of 1,140k generative manipulated images with their corresponding origin images. Moreover, the dataset can be split into subsets. GIM-SD-all utilizes the training set from Imagenet-1k [10] and employs the Stable Diffusion [53] generator for image manipulation dedicated to manipulation research. Through the analysis of the data scale, GIM-SD, GIM-GLIDE and GIM-DDNM individually select a random set of 100 classes from the Imagenet-1k training dataset and the entire test data to generate their respective training and test sets to benchmark the image manipulation detection and location (IMDL)methods. GIM-SD (VOC) is a cross-data-distribution test set constructed using the test set from the PASCAL VOC [14] dataset and Stable Diffusion techniques.
@@
The GIM data format consists of JPG images paired with PNG labels. Filenames ending with "_f" indicate tampered images. The masks segmented by SAM [33] serve as ground truth for manipulation detection, containing two labels denoted by 0 and 255, representing the original and manipulated categories, respectively.
6.2 Dataset Comparison.
GIM focuses on generative local manipulation in natural images and provides manipulated region labels, specifically for image manipulation detection and localization tasks. This distinguishes GIM from the existing datasets focused on faces or limited to detection (classification). As shown in Table 10, we compare several local image manipulation datasets. The existing image forensic datasets are primarily constructed using traditional manipulation techniques. These datasets commonly employ manual or random manipulation to generate data. AutoSplice [29], HiFi-IFDL[20], CoCoGlide  [19] dataset utilizes generative algorithms for image manipulation. However, they rely on one type of manipulation and suffer from a small data volume.
@@
Columbia Gray [43] 933 912
Columbia Color [24] 183 180
MICC-F2000 [1] 1,300 700
VIPP Synth.  [1] 4,800 4,800
CASIA V1.0  [13] 800 921
CASIA V2.0  [13] 7,200 5,123
Wild Web [66] 90 9,657
NC2016  [17] 560 564
NC2017  [17] 2,667 1,410
MFC2018  [17] 14,156 3,265
MFC2019  [17] 10,279 5,750
PS-Battles [22] 11,142 102,028
DEFACTO  [42] 0 229,000
IMD2020 [45] 35,000 35,000
SP COCO  [36] 0 200,000
CM COCO  [36] 0 200,000
CM RAISE  [36] 0 200,000
HiFi-IFDL[20] - 1,000,00*
CoCoGlide [19] 0 512
AutoSplice  [29] 3,621 2,273
@@
7 Implementation Details
7.1 Architecture
@@ 7.2 Training Process
For the training of the ShadowTracer network $g_{\phi}$, We randomly selected 20,000 authentic images from ImageNet and correspondingly generated manipulated images. During comprehensive performance verification, all three generators (Stable Diffusion [52], GLIDE [44], DDNM [59])are used, whereas for generalization verification, only Stable Diffusion is employed. Throughout the training, the alpha blending parameter ranges between 0.5 and 1.0, randomly producing blended images subjected to three degradation types (JPEG compression, downsampling, Gaussian blur). The network is trained on 40$\times$40 pixel patches randomly sampled from the dataset. Training is conducted for roughly 300,000 iterations with a batch size of 64. An Adam [31] optimizer is employed, initialized with a learning rate of 0.001.
@@
8 Additional Ablation Analysis
@@
9 More Visualizations on GIM
In this section, we present additional visualizations of the GIM dataset in Figure  7,  8, and  9, showcasing the data visualization from GIM-SD, GIM-GLIDE, and GIM-DDNM respectively.
10 Societal Impact
Our research yields a positive societal impact on the community by focusing on addressing the challenge of detecting and locating generative-based manipulations. We introduce a reliable database, GIM, aimed at enhancing the security of AI-generated content (AIGC). This database facilitates the training of Image Manipulation Detection and Localization (IMDL)methods on GIM, allowing them to generalize across various scenarios. Consequently, our algorithm GIMFormer fosters increased trust among the general public in our society regarding media content. GIM and GIMFormer are beneficial for digital media forensics, especially generative manipulation with real-world degradations.
11 Ethics Statement
Our GIM is based on ImageNet and VOC. No additional personally identifiable information or sensitive personally identifiable information is introduced during the production of fake images in the GIM dataset. During the dataset production, we do not introduce extra information containing exacerbated bias against people of a certain gender, race, sexuality, or who have other protected characteristics. The ethical issues in the ImageNet and VOC datasets have been discussed in previous works. Crawford et al.  [8] discuss issues with ImageNet. The first issue is the political nature of all taxonomies or classification systems, where terms like "male" and "female" are considered "natural," while "hermaphrodite" is offensively placed within the branch of Person > Sensualist > Bisexual alongside "pseudohermaphrodite" and "switch hitter" categories. The second issue concerns offensive images of real people, while the third is the use of people’s photos without their consent by ImageNet creators.

### v2

@@ Appendix A GIM Dataset
The specific quantities and details of the dataset are presented in Table 9, GIM has a total number of 1,140k generative manipulated images with their corresponding origin images. Moreover, the dataset can be split into subsets. GIM-SD-all utilizes the training set from Imagenet-1k (Deng et al. 2009) and employs the Stable Diffusion generator for image manipulation dedicated to manipulation research. Through the analysis of the data scale, GIM-SD, GIM-GLIDE and GIM-DDNM individually select a random set of 100 classes from the Imagenet-1k training dataset and the entire test data to generate their respective training and test sets to benchmark the image manipulation detection and location (IMDL)methods. GIM-VOC is a cross-data-distribution test set constructed using the test set from the PASCAL VOC (Everingham et al. 2010) dataset and Stable Diffusion techniques.
@@
The GIM data format consists of JPG images paired with PNG labels. Filenames ending with ”_f” indicate tampered images. The masks segmented by SAM (Kirillov et al. 2023b) serve as ground truth for manipulation detection, containing two labels denoted by 0 and 255, representing the original and manipulated categories, respectively.
Dataset Comparison.
GIM focuses on generative local manipulation in natural images and provides manipulated region labels, specifically for image manipulation detection and localization tasks. This distinguishes GIM from the existing datasets focused on faces or limited to detection (classification). As shown in Table 10, we compare several local image manipulation datasets. The existing image forensic datasets are primarily constructed using traditional manipulation techniques. These datasets commonly employ manual or random manipulation to generate data. AutoSplice (Jia et al. 2023), HiFi-IFDL(Guo et al. 2023), CoCoGlide  (Guillaro et al. 2023) dataset utilizes generative algorithms for image manipulation. However, they rely on one type of manipulation and suffer from a small data volume.
@@
Columbia Gray (Ng, Chang, and Sun 2004) 933 912
Columbia Color (Hsu and Chang 2006) 183 180
MICC-F2000 (Amerini et al. 2011) 1,300 700
VIPP Synth.  (Amerini et al. 2011) 4,800 4,800
CASIA V1.0  (Dong, Wang, and Tan 2013) 800 921
CASIA V2.0  (Dong, Wang, and Tan 2013) 7,200 5,123
Wild Web (Zampoglou, Papadopoulos, and Kompatsiaris 2015) 90 9,657
NC2016  (Guan et al. 2019) 560 564
NC2017  (Guan et al. 2019) 2,667 1,410
MFC2018  (Guan et al. 2019) 14,156 3,265
MFC2019  (Guan et al. 2019) 10,279 5,750
PS-Battles (Heller, Rossetto, and Schuldt 2018) 11,142 102,028
DEFACTO  (Mahfoudi et al. 2019) 0 229,000
IMD2020 (Novozamsky, Mahdian, and Saic 2020) 35,000 35,000
SP COCO  (Kwon et al. 2022) 0 200,000
CM COCO  (Kwon et al. 2022) 0 200,000
CM RAISE  (Kwon et al. 2022) 0 200,000
HiFi-IFDL(Guo et al. 2023) - 1,000,00*
CoCoGlide (Guillaro et al. 2023) 0 512
AutoSplice  (Jia et al. 2023) 3,621 2,273
@@
Appendix B Implementation Details
Architecture
@@ Appendix B Implementation Details
For the training of the ShadowTracer network $g_{\phi}$, We randomly selected 20,000 authentic images from ImageNet and correspondingly generated manipulated images. During comprehensive performance verification, all three generators (Stable Diffusion (Rombach et al. 2022), GLIDE (Nichol et al. 2021), DDNM (Wang, Yu, and Zhang 2022))are used, whereas for generalization verification, only Stable Diffusion is employed. Throughout the training, the alpha blending parameter ranges between 0.5 and 1.0, randomly producing blended images subjected to three degradation types (JPEG compression, downsampling, Gaussian blur). The network is trained on 40$\times$40 pixel patches randomly sampled from the dataset. Training is conducted for roughly 300,000 iterations with a batch size of 64. An Adam (Kingma and Ba 2014) optimizer is employed, initialized with a learning rate of 0.001.
@@
Appendix C Additional Ablation Analysis
@@
Appendix D More Visualizations on GIM
In this section, we present additional visualizations of the GIM dataset in Figure  7,  8, and  9, showcasing the data visualization from GIM-SD, GIM-GLIDE, and GIM-DDNM respectively. These figures illustrate the diverse data representations within the GIM dataset, showcasing the unique characteristics brought by different generative models and manipulation techniques.
@@ Appendix D More Visualizations on GIM
Appendix E Societal Impact
Our research contributes positively to society by addressing the challenge of detecting and locating generative-based manipulations, ultimately enhancing the security of AI-generated content (AIGC). The introduction of the reliable GIM database aims to bolster the effectiveness of Image Manipulation Detection and Localization (IMDL) methods, allowing for generalization across diverse scenarios. As a result, our algorithm GIMFormer instills greater trust among the general public in our society regarding media content authenticity. The utilization of GIM and GIMFormer proves advantageous for digital media forensics, particularly in addressing generative manipulation with real-world degradations. The practical applications of the GIM dataset have the potential to create a significant social impact. For instance, the training of detectors on GIM enables users to identify AI-generated content, leading to consequential actions such as identifying academic dishonesty. Widely disseminating and utilizing the GIM dataset for IMDL algorithms training could pave the way for its integration into specialized software designed for this purpose, thereby enhancing the accuracy and reliability of algorithms. However, it is essential to note that models trained on this dataset may exhibit undesirable tendencies towards novel objects in the real world, necessitating careful consideration for practical implementation despite their demonstrated generalizability.
Appendix F Ethics Statement
Our GIM is based on ImageNet and VOC. No additional personally identifiable information or sensitive personally identifiable information is introduced during the production of fake images in the GIM dataset. During the dataset production, we do not introduce extra information containing exacerbated bias against people of a certain gender, race, sexuality, or who have other protected characteristics. The ethical issues in the ImageNet and VOC datasets have been discussed in previous works. Crawford et al.  (Crawford and Paglen 2021) discuss issues with ImageNet. The first issue is the political nature of all taxonomies or classification systems, where terms like ”male” and ”female” are considered ”natural,” while ”hermaphrodite” is offensively placed within the branch of Person ¿ Sensualist ¿ Bisexual alongside ”pseudohermaphrodite” and ”switch hitter” categories. The second issue concerns offensive images of real people, while the third is the use of people’s photos without their consent by ImageNet creators.
Appendix G Limitations
The GIM dataset shares classes with ImageNet and VOC, but it may not encompass future emerging objects due to the evolving variety in the real world. Fine-tuning pre-trained models on new object data can address this issue. Current research primarily centers on image manipulation. The emergence of realistic video generation models like SORA presents fresh challenges as AI manipulation extends into videos. Future plans involve expanding research to address the complexities of video manipulation.
Appendix H Acknowledgments
This work is partially supported by NSFC ( No. 62402318, 62376153, 24Z990200676)